OpenStudy (anonymous):

There are x benches and y students in an auditorium. When a bench is seated with 3 students, 18 are left standing. If a bench is seated with 4 students, 3 benches would be empty. a. Write 2 equations showing the relationship between x benches and y students b. Find the number of benches and students

OpenStudy (anonymous):

x- benches, y-students y=3x+18 x-3=y/4

OpenStudy (anonymous):

Thanks, what about b?

OpenStudy (anonymous):

you need to try to solve it by yourself :-) I will help you if you will need it.

OpenStudy (anonymous):

Well alright, no prob there, thanks for the help again

OpenStudy (anonymous):

you are welcome. Do you understand it?

OpenStudy (anonymous):

Kind of, think I could figure it out by myself though

OpenStudy (anonymous):

y=3x+18 y=4(x-3) right parts are equal and you can find x

OpenStudy (anonymous):

Yep, I get it

OpenStudy (anonymous):

answer is?

OpenStudy (anonymous):

one sec

OpenStudy (anonymous):

x=30 right?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

so y=108

OpenStudy (anonymous):

yes, you are right. With system of equations always good to do check for both equations.

OpenStudy (anonymous):

Sweet

OpenStudy (anonymous):

glad to help

OpenStudy (anonymous):

x- benches, y-students y=3x+18 x-3=y/4 y=3x+18 y=4(x-3) x = 30 y = 108